OpenStudy (dinorap1):

PLEASE HELP! WILL GIVE MEDAL TO BEST ANSWER. @mathmate Solve the system using substitution. {-x-y-z=-8 {-4x+4y+5z=7 {2x+2z=4

OpenStudy (dinorap1):

Nevermind, @mathmate isn't online. Anyone else want to help???

OpenStudy (anonymous):

may i?

OpenStudy (dinorap1):

Yes, of course! :D

OpenStudy (dinorap1):

The solution set I got was (2,7,-1) and it's wrong, so I'm getting really frustrated right now...

OpenStudy (anonymous):

is it must use substitution only?

OpenStudy (dinorap1):

Yep.

OpenStudy (dinorap1):

Well, I'm supposed to show my work, so if I do it another way it'll probably be marked wrong...

OpenStudy (dinorap1):

Oh hello, @mathmate ! You helped me before so I thought you could help me with another one!

OpenStudy (mathmate):

Examine the equations and decide which is the best one to use for substitution!

OpenStudy (dinorap1):

The first equation is, I'm pretty sure. I reduced it down to z=8-y-x

OpenStudy (mathmate):

You are substituting one variable with two other ones. What if you take (3)/2 which gives x+z=2 Transpose the x to the right gives z=2-x .......(4) Substitute z (from (4)) into (1) and (2) and the trick is done. You end up with two equations with unknowns x and z.

OpenStudy (anonymous):

-y+8=x+z |2 -2y+16= 2x+2z

OpenStudy (mathmate):

*x and y

OpenStudy (dinorap1):

Wait... so you're dividing the 3rd equation by 2 @mathmate ?

OpenStudy (mathmate):

yes, there is a common factor of two.

OpenStudy (dinorap1):

Hmmm... that's a different way to do it! Never would've thought of doing it that way! I'll try it your way now and see what happens...

OpenStudy (mathmate):

What you suggested is ok too! But it might be a little more arithmetic operations.

OpenStudy (dinorap1):

When I tried it my way, I got mostly decimals. So I got really annoyed with the equation I chose... So when I plug (z=2-x) into -x-y-z=-8, is there any distributing involved?

OpenStudy (mathmate):

First, i would change -x-y-z=-8 to x+y+z=8 (life is easier that way). Then x+y+(2-x)=8 That gives y=6 right away! lol

OpenStudy (dinorap1):

But are you really allowed to get rid of all the negative signs like that? o_o

OpenStudy (mathmate):

Yes, basically you multiply both sides by -1.

OpenStudy (dinorap1):

Wow, okay! I wish I knew as many little tricks like that! Okay, so now how do I figure out the other 2 variables? I forgot...

OpenStudy (mathmate):

Call y=6 ..............(5) Substitute (4) and (5) in (2) to find x. Substitute this value of x into (4) to find z.

OpenStudy (dinorap1):

Okay, I'll do that and see what I get...

OpenStudy (mathmate):

THe best trick is to write down the equations (and the numbers like .......(4)) neatly in order, then you'll find your way.

OpenStudy (dinorap1):

I got: (3, 6, -1)

OpenStudy (mathmate):

That's perfect. Another trick is to substitute what you've got back into equations (1), (2) and (3) and check that they all work. This is an exam saver! :)

OpenStudy (dinorap1):

Yep, I just checked, and it's right! Yay, thanks for helping me solve another one! You're the best! :D Could you help me with another couple of questions...?

OpenStudy (mathmate):

Sorry, I have to go. However, I love to give you a medal because for the longest time I haven't worked with someone who is actually interested in working out the problems! You're the best!

OpenStudy (dinorap1):

Aw, alright! I guess I'll have to post another question on here and wait for someone to help... Aww, thanks! I appreciate it! I'll write a testimonial for you once I get the chance! :D

OpenStudy (dinorap1):

Oops, I didn't mean to post that screenshot. That's the problem I was gonna ask for help for. Just ignore that! :)